A thief in stolen car crashes through a police checkpost at top speed of 90 kmph. A motorcycle cop, reacting after 2 sec, accelerates from rest at 5m/s2.His top spd is 108 kmph.


Find max. separation between theif and police?


a)112.5m b)115m c)116.5m d)None above


Also plz. explain me how to solve these kind of problems when we have to find separation, top speed etc.

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear chilukuri

speed of the car = 90 kmph = 25 m/sec

let after t sec of police stated distance between  them are maximum

max speed of police =108 km/h = 30 m/sec

time taken by police to atttain a speed of 25 m/sec=25/5 =5 sec

distance between them will increse  till police man attain a speed og more than 25 m/sec

 distance travell be thief = 25*(5+2) =175 m

distance travell by police =1/2 *5* 5=62.5 m

so distance between them S=175 -62.5 =112.5


 

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin